# Fractal Patterns May Unravel the Intelligence in Next-Token Prediction

### v1

@@ Appendix C Predicting Downstream Performance
Please note that our results are not directly comparable to all previous published results for the same models; please cite the original results from (Chowdhery et al., 2022; Anil et al., 2023b). Here, we only aim for a fair comparison between models: only pretrained models without instruction tuning are used, we do not optimize any prompts for each model, and we evaluate all models using only a 2K sequence length.
Model
BPB
0S BBH Direct
0S BBH CoT
0S MMLU
3S BBH Direct
3S BBH CoT
5S MMLU
8S GSM8K CoT
0S BBH +MMLU
FS BBH +MMLU +GSM8K
T5-Decoder-110M
1.11
0.83
0.11
25.65
21.36
5.69
25.62
0.91
13.06
13.35
T5-Decoder-340M
1.00
0.96
0.17
25.72
23.57
10.03
25.98
1.59
13.14
14.79
T5-Decoder-1B
0.92
1.29
0.14
25.99
24.26
13.19
24.82
1.14
13.35
14.90
T5-Decoder-5B
0.85
2.13
0.48
24.41
24.76
18.05
25.63
2.20
12.86
16.41
PaLM1-8B
0.78
6.46
1.21
23.53
32.18
27.60
24.56
5.16
13.68
19.87
PaLM1-62B
0.70
13.79
0.83
51.86
39.51
39.70
54.78
29.57
29.59
41.32
PaLM1-540B
0.66
23.26
4.72
67.78
52.44
56.02
70.50
56.79
40.89
60.51
PaLM2-XXS
0.81
8.99
0.13
25.26
30.71
26.08
24.72
2.96
14.91
18.69
PaLM2-XS
0.73
16.68
0.95
49.69
38.28
37.64
47.42
22.14
29.25
35.84
PaLM2-S
0.67
23.60
4.24
69.89
48.88
50.88
68.12
50.49
41.91
56.16
PaLM2-M
0.65
21.32
5.70
69.62
52.49
56.04
69.33
59.21
41.57
60.94
PaLM2-L
0.61
24.00
10.19
79.10
66.34
66.66
78.64
80.36
48.10
75.17
@@
BPB
$\mathrm{S}$
$\mathrm{J}$
BPB+$\mathrm{S}$
BPB+$\mathrm{J}$
0S BBH Direct
0.785
-0.060
0.673
0.761
0.794
0S MMLU
0.653
-0.067
0.426
0.614
0.614
0S BBH+MMLU
0.685
-0.065
0.472
0.650
0.651
3S BBH Direct
0.767
-0.030
0.599
0.744
0.754
3S BBH CoT
0.881
-0.026
0.678
0.870
0.879
5S MMLU
0.660
-0.044
0.421
0.624
0.622
8S GSM8K CoT
0.654
-0.037
0.427
0.619
0.616
FS BBH + MMLU+GSM8K
0.717
-0.036
0.489
0.687
0.686

### v2

@@ Appendix C Predicting Downstream Performance
Please note that our results are not directly comparable to all previous published results for the same models; please cite the original results from [12, 6]. Here, we only aim for a fair comparison between models: only pretrained models without instruction tuning are used, we do not optimize any prompts for each model, and we evaluate all models using only a 2K sequence length.
Model BPB 0S BBH Direct 0S BBH CoT 0S MMLU 3S BBH Direct 3S BBH CoT 5S MMLU 8S GSM8K CoT 0S BBH +MMLU FS BBH +MMLU +GSM8K
T5-Decoder-110M 1.11 0.83 0.11 25.65 21.36 5.69 25.62 0.91 13.06 13.35
T5-Decoder-340M 1.00 0.96 0.17 25.72 23.57 10.03 25.98 1.59 13.14 14.79
T5-Decoder-1B 0.92 1.29 0.14 25.99 24.26 13.19 24.82 1.14 13.35 14.90
T5-Decoder-5B 0.85 2.13 0.48 24.41 24.76 18.05 25.63 2.20 12.86 16.41
PaLM1-8B 0.78 6.46 1.21 23.53 32.18 27.60 24.56 5.16 13.68 19.87
PaLM1-62B 0.70 13.79 0.83 51.86 39.51 39.70 54.78 29.57 29.59 41.32
PaLM1-540B 0.66 23.26 4.72 67.78 52.44 56.02 70.50 56.79 40.89 60.51
PaLM2-XXS 0.81 8.99 0.13 25.26 30.71 26.08 24.72 2.96 14.91 18.69
PaLM2-XS 0.73 16.68 0.95 49.69 38.28 37.64 47.42 22.14 29.25 35.84
PaLM2-S 0.67 23.60 4.24 69.89 48.88 50.88 68.12 50.49 41.91 56.16
PaLM2-M 0.65 21.32 5.70 69.62 52.49 56.04 69.33 59.21 41.57 60.94
PaLM2-L 0.61 24.00 10.19 79.10 66.34 66.66 78.64 80.36 48.10 75.17
@@
BPB $\mathrm{S}$ $\mathrm{J}$ BPB+$\mathrm{S}$ BPB+$\mathrm{J}$
0S BBH Direct 0.785 -0.060 0.673 0.761 0.794
0S MMLU 0.653 -0.067 0.426 0.614 0.614
0S BBH+MMLU 0.685 -0.065 0.472 0.650 0.651
3S BBH Direct 0.767 -0.030 0.599 0.744 0.754
3S BBH CoT 0.881 -0.026 0.678 0.870 0.879
5S MMLU 0.660 -0.044 0.421 0.624 0.622
8S GSM8K CoT 0.654 -0.037 0.427 0.619 0.616
FS BBH + MMLU+GSM8K 0.717 -0.036 0.489 0.687 0.686